# Temporal and Semantic Evaluation Metrics for Foundation Models in Post-Hoc Analysis of Robotic Sub-tasks

### v1

@@ V Experiments
The goal of our experiments is to verify our approach on a large and diverse set of trajectory lengths, initial conditions, sub-task durations, and sub-task descriptions. Our experiments include 200 unique trajectories (4 environments, 50 trajectories per environment) with corresponding $\mathcal{D}$ and a ground-truth $\mathcal{S}$. All project code including experimental configurations, Algorithm 1 implementation, FSM implementations, and results generation can be found at github.com/jsalfity/task_decomposition.
@@ V-A Robotics’ Trajectory Data Collection
We narrow the scope of our work to robot manipulators in the Robosuite [42] simulation environment. The 4 different environments used in this work are seen in Table II including their approximate $K_{\mathcal{S}}$ and $N_{\mathcal{S}}$. The environments range from short, single objective tasks to longer, complicated multi-objective tasks. All environments are controlled using FSM-based control, seen in the supporting code, to control robot’s end-effector position based on pre-defined sub-tasks. The transition between sub-tasks is determined by spatial position conditions such as when a gripper is directly above an object, when a grasp is achieved, etc. Each sub-task is labeled directly by the FSM. As an example, the Lift environment’s sequential sub-task descriptors are: “Move to cube”, “Grasp Cube”, “Lift Cube”, and moves to the next sub-task when the previous sub-task is complete. The FSM approach enables tracking specific sub-tasks and the associated start/stop steps. In each sub-task, a position feedback controller commands the end-effector to move to the desired x-y-z position using Robosuite’s OSC_POSE controller.
@@
To generate a unique trajectory, environmental objects are randomly positioned. Thus the general manipulation task for each environment remain similar but each sub-task contains different start and end steps and different $K_{\mathcal{S}}$. The $N_{\mathcal{S}}$ and approximate $K_{\mathcal{S}}$ is written below each environment in Table II.
@@ V-B Querying an FM
Four FM models are used: GPT-4V [8], Gemini Pro, Gemini Pro Vision [9], and Video-LLaVA [10]. The configuration settings for each model can be found in Section VII and the supporting code. These models offer a diversity of input modality (textual and visual) and different parametric tuning accessibility and customization. For example, Video-LLaVA is an open source model, whereas GPT-4V is only accessed by an API.
For each FM query, the P is built how it appears in Fig. 2. The only differences are $\textbf{C}^{1S}$ for each environment and $\mathcal{D}$ unique to each trajectory. The same $\textbf{C}^{1S}$ (from a hold-out $\mathcal{D}$, $\mathcal{S}$ pair) is used for each environment and consists of important transitions between sequential sub-task decompositions.
V-C Results
The experimental results, shown in Table II, are ablated across different inputs: One-Shot context, $\textbf{C}^{1S}$, Zero-Shot context, $\textbf{C}^{0S}$, textual data, $\mathcal{D}^{kxu}$, visual data, $\mathcal{D}^{\mu}$, and both textual and visual data, $\mathcal{D}$. Gemini Pro Vision and Video-LLaVa could not accept $\mathcal{D}$ due to their respective P token limit and therefore do not have such statistics.
Statistics from comparing ground-truth against a random sub-task decomposition, $\tilde{\mathcal{S}}$, is included to provide a baseline comparison. An example is shown at the right-most column of Table I and statistics are in Table II and Fig. 3. The randomized baseline for each environment is formed by $\texttt{SIMILARITY}(\mathcal{S},\tilde{\mathcal{S}})$. Each $\tilde{\mathcal{S}}$ is synthesized from a function that generates a sub-task decomposition with total length $K_{\mathcal{S}}$, but results in $N_{\tilde{\mathcal{S}}}$ random sub-tasks. Each $\tilde{s}_{n}$ is defined by a uniformly random chosen start and end step and a randomly chosen description from POSSIBLE_SUBTASKS. The function ensures there is no overlap in steps between sub-tasks. In the majority of cases, $N_{\tilde{\mathcal{S}}}\neq N_{\mathcal{S}}$, which practically means the granularity of decomposition is different. POSSIBLE_SUBTASKS contains 20 distinct $\zeta$ that are all candidates for a manipulator task to achieve. The exact function and POSSIBLE_SUBTASKS list is shown in the github repository.
Table II shows statistics over FM responses with $\hat{\mathcal{S}}$ that are valid according to Definition 3. Examples of invalid FM responses include when $k_{\text{start}}$ comes after $k_{\text{end}}$ for any given $\hat{s}_{n}$, $k_{\text{start}}$ in any $\hat{s}_{n}$ is greater than $k_{\text{start}}$ in any $\hat{s}_{n+1}$, or the FM response contains no $\hat{\mathcal{S}}$ at all. The error rates for each batch of 50 FM calls, due to 50 trajectories per environment, is reported in Table III.
It is clear from the results in Table II and the two environment examples in Fig. 3 that there is significant advantage when using $\textbf{C}^{1S}$, as all $\tau_{k}$ and $\tau_{\zeta}$ are higher compared to using $\textbf{C}^{0S}$. We speculate that this is because the in-context example directly steers the FM to attend to important transitional steps and to synthesize descriptors using similar terminology to the example.
The very low scores are consistent for $\hat{\mathcal{S}}$ that contain $N_{\hat{\mathcal{S}}}\gg N_{\mathcal{S}}$, indicating too fine of a granularity. For example, Video-LLaVa produced $N_{\hat{\mathcal{S}}}\approx K_{\mathcal{S}}$, implying a sub-task at every $k$, which is not realistic.
@@ V-C Results
LLM (Context, Data) $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$
GPT-4V $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ $0.97\pm 0.07$ $0.87\pm 0.16$ $0.80\pm 0.15$ $0.83\pm 0.10$ $0.93\pm 0.10$ $0.60\pm 0.12$ $0.74\pm 0.12$ $0.90\pm 0.08$
@@
Gemini Pro $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ $0.92\pm 0.06$ $0.98\pm 0.02$ $0.84\pm 0.15$ $0.52\pm 0.21$ $0.62\pm 0.13$ $0.65\pm 0.14$ $0.53\pm 0.15$ $0.88\pm 0.05$
@@
Gemini Pro V $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.91\pm 0.07$ $0.84\pm 0.15$ $0.61\pm 0.18$ $0.74\pm 0.24$ $0.57\pm 0.19$ $0.67\pm 0.14$ $0.54\pm 0.17$ $0.89\pm 0.06$
@@
Video-LLaVA $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.91\pm 0.06$ $0.98\pm 0.01$ $0.58\pm 0.11$ $0.65\pm 0.02$ $0.64\pm 0.13$ $0.67\pm 0.14$ $0.32\pm 0.17$ $0.54\pm 0.01$
@@
Random Baseline $0.33\pm 0.09$ $0.21\pm 0.10$ $0.43\pm 0.13$ $0.24\pm 0.04$ $0.36\pm 0.15$ $0.20\pm 0.06$ $0.32\pm 0.10$ $0.21\pm 0.06$
TABLE II: Ablation study across all environments: Comparison of temporal similarity, $\tau_{k}$, and semantic similarity, $\tau_{\zeta}$, for different robotics manipulation environments ablated across prompts, P, with different inputs modalities (textual data, $\mathcal{D}^{kxu}$ and visual data, $\mathcal{D}^{\mu}$) and context, C, (in-context example, $\textbf{C}^{1S}$, and no in-context example, $\textbf{C}^{0S}$). Statistics (mean and standard deviation) are reported over approximately 50 predictions for each environment, FM, input modality, and context combinations. The key insight is the significant increase in temporal and semantic similarities for $\hat{\mathcal{S}}$ with P that include $\mathcal{D}^{kxu}$ and $\textbf{C}^{1S}$.
@@ VI Conclusion
We present a method to send multi-modal input data including step-by-step state, action and corresponding images to FM including LLMs and VLMs. We steer an FM for use as an analytics engine to digest raw trajectory data in the prompt and query the FM to decompose the data into distinct sub-tasks. As of this writing, the cost of offloading and requesting data is \$0.03 to \$0.15 depending on the trajectory length. We see this analysis and capability of using FMs for robotic’s data labeling as cost-effective and data efficient for the broader research problem of robot learning at scale.
The results are shown with a ground-truth $\mathcal{S}$ crafted by a customized FSM. There may be other $\mathcal{S}$ with different granularity of sub-tasks, resulting in a different number of sub-tasks, $N_{\mathcal{S}}$, and alternative descriptors. But our results are still valid because the proposed annotation pipeline and evaluation method will be set up by developers who want FM predictions, $\hat{\mathcal{S}}$, and ensuing annotations similar to $\mathcal{S}$ of their own.
The IOU calculation is a strict temporal calculation, resulting in strict weightings. This could create challenges for future developers, but an alternative “soft” IOU could be utilized where an intersection may be greater than zero for temporal ranges that are near each other, but do not necessarily intersect. In our algorithm, shorter sub-tasks are assigned smaller weights towards $\tau_{k}$ and $\tau_{\zeta}$. In practice, a short sub-task may be more important to a developer than longer sub-tasks. Users can manually increase the relative weighting of important sub-tasks to give more contribution towards the final scores. While we use the USE encoder to calculate $\tau_{\zeta}$, our framework is designed to be modular, allowing future use of any pre-trained encoder.
The presented robotics manipulation environments have structured and repeatable $\mathcal{S}$, which then allow us to present quantitative results. Due to limits on prompt tokens, longer trajectories - given the resources - can also be considered using our approach.

### v2

@@ V Experiments
The goal of our experiments is to verify our approach on a large and diverse set of trajectory lengths, initial conditions, sub-task durations, and sub-task descriptions. Our experiments include 200 unique trajectories (4 environments, 50 trajectories per environment) with corresponding $\mathcal{D}$ and a ground-truth $\mathcal{S}$. All project code including experimental configurations, Algorithm 1 implementation, FSM implementations, and results generation can be found in the github repository github.com/jsalfity/task_decomposition.
@@ V-A Robotics’ Trajectory Data Collection
Robotics’ Environments: We narrow the scope of our work to robot manipulators in the Robosuite [36] simulation environment. While the following experiments and results are limited to Robosuite, we note that any simulation environment or real-world experimental setup which can produce trajectory data, $\mathcal{D}$, in the form defined in Definition 1 can be similarly evaluated. The 4 different environments used in this work are seen in Table II including their approximate $K_{\mathcal{S}}$ and $N_{\mathcal{S}}$. The environments range from short, single objective tasks to longer, complicated multi-objective tasks.
Finite-State Machine: All environments are controlled using FSM-based control, seen in the supporting code, to control robot’s end-effector position based on pre-defined sub-tasks. The transition between sub-tasks is determined by spatial position conditions such as when a gripper is directly above an object, when a grasp is achieved, etc. Each sub-task is labeled directly by the FSM. As an example, the Lift environment’s sequential sub-task descriptors are: “Move to cube”, “Grasp Cube”, “Lift Cube”, and moves to the next sub-task when the previous sub-task is complete. The FSM approach enables tracking specific sub-tasks and the associated start/stop steps. In each sub-task, a position feedback controller commands the end-effector to move to the desired x-y-z position using Robosuite’s OSC_POSE controller.
@@
Environmental objects are randomly positioned to generate unique trajectories. The general manipulation task for each environment remain similar but each sub-task contains different start and end steps and different $K_{\mathcal{S}}$. The $N_{\mathcal{S}}$ and approximate $K_{\mathcal{S}}$ are written below each environment in Table II.
Ablation: Although all this information is available to us through the simulated environment, we conduct an ablation study in our experiments, Fig. 3, to vary the amount of information provided to each FM for sub-task decomposition. Notably, $\mathcal{D}^{\mu}$, which represents scenarios where only RGB camera data is available for labeling, best reflects real-world use cases.
Human Annotations: $\texttt{SIMILARITY}(\mathcal{S},\tilde{\mathcal{S}})$ statistics with ground-truth, $\mathcal{S}$, and human annotated sub-task decompositions, $\tilde{\mathcal{S}}$, provide a baseline comparison. Human annotators in our experiments constitute roboticists from the authors’ academic group and are not biased by the authors. Human annotations are collected by presenting annotators with only visual data, $\mathcal{D}^{\mu}$, i.e. the video, and similar instructions as shown in Fig. 2. A partial in-context example is shown for each video, which guides the $\tilde{\mathcal{S}}$ granularity and vernacular. As each human annotation is timed, we learned constructing a $\tilde{\mathcal{S}}$ takes 2-3 minutes, which is nearly 5x the cost of an FM-based annotation. An example $\tilde{\mathcal{S}}$ is shown at the right-most column of Table I and statistics are in Table II and Fig. 3.
@@ V-B Querying an FM
Four FM models are used: GPT-4V [9], Gemini Pro, Gemini Pro Vision [10], and Video-LLaVA [11]. The configuration settings for each model can be found in Section VII and the supporting code. These models offer a diversity of input modality (textual and visual) and different parametric tuning accessibility and customization, e.g., Video-LLaVA is open source, whereas GPT-4V is only accessible via API.
For each FM query, the P is constructed as shown in Fig. 2, with variations only in $\textbf{C}^{1S}$ for each environment and $\mathcal{D}$ for each trajectory. The same $\textbf{C}^{1S}$ is used across environments and is based on a hold-out pair of $\mathcal{D}$ and $\mathcal{S}$, capturing key transitions between sequential sub-task decompositions.
Statistics: Table II shows statistics over FM responses with $\hat{\mathcal{S}}$ that are valid according to Definition 3. The error rates for each batch of 50 FM calls, due to 50 trajectories per environment, is reported in Table III, which is the corresponding sample size for the mean and standard deviation of each $\tau_{k},\tau_{\zeta}$ pair in Table II. Examples of invalid FM responses include when $k_{\text{start}}$ comes after $k_{\text{end}}$, $k_{\text{start}}$ in any $\hat{s}_{n}$ is greater than $k_{\text{start}}$ in any $\hat{s}_{n+1}$, or the FM response contains no $\hat{\mathcal{S}}$ at all. Gemini Pro can only take in textual inputs; Gemini Pro Vision and Video-LLaVa could not accept $\mathcal{D}$ due to their respective P token limit and therefore only have $\mathcal{D}^{\mu}$ statistics.
@@
FM (Context, Data) $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$ $\tau_{k}$ $\tau_{\zeta}$
Human Annotations $(\mathcal{D}^{\mu})$ $0.64\pm 0.11$ $0.48\pm 0.12$ $0.81\pm 0.12$ $0.45\pm 0.14$ $0.63\pm 0.14$ $0.41\pm 0.14$ $0.66\pm 0.16$ $0.38\pm 0.13$
GPT-4V $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ $\textbf{0.97}\pm\textbf{0.07}$ $0.87\pm 0.16$ $0.80\pm 0.15$ $\textbf{0.83}\pm\textbf{0.10}$ $\textbf{0.93}\pm\textbf{0.10}$ $0.60\pm 0.12$ $\textbf{0.74}\pm\textbf{0.12}$ $\textbf{0.90}\pm\textbf{0.08}$
@@
Gemini Pro $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ $0.92\pm 0.06$ $\textbf{0.98}\pm\textbf{0.02}$ $\textbf{0.84}\pm\textbf{0.15}$ $0.52\pm 0.21$ $0.62\pm 0.13$ $0.65\pm 0.14$ $0.53\pm 0.15$ $0.88\pm 0.05$
@@
Gemini Pro V $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.91\pm 0.07$ $0.84\pm 0.15$ $0.61\pm 0.18$ $0.74\pm 0.24$ $0.57\pm 0.19$ $\textbf{0.67}\pm\textbf{0.14}$ $0.54\pm 0.17$ $0.89\pm 0.06$
@@
Video-LLaVA $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ $0.91\pm 0.06$ $\textbf{0.98}\pm\textbf{0.01}$ $0.58\pm 0.11$ $0.65\pm 0.02$ $0.64\pm 0.13$ $\textbf{0.67}\pm\textbf{0.14}$ $0.32\pm 0.17$ $0.54\pm 0.01$
@@
TABLE II: Ablation study across all environments: Comparison of temporal similarity, $\tau_{k}$, and semantic similarity, $\tau_{\zeta}$, for different robotics manipulation environments ablated across prompts, P, with different inputs modalities (textual data, $\mathcal{D}^{kxu}$ and visual data, $\mathcal{D}^{\mu}$) and context, C, (in-context example, $\textbf{C}^{1S}$, and no in-context example, $\textbf{C}^{0S}$). Statistics (mean and standard deviation) are reported over approximately 50 predictions for each environment, FM, input modality, and context combinations. The key insight is the significant increase in temporal and semantic similarities for $\hat{\mathcal{S}}$ with P that include $\mathcal{D}^{kxu}$ and $\textbf{C}^{1S}$. Statistics are in bold highlight exemplary high values for each metric and environment.
V-C Results
Our experimental results, seen completely in Table II, show that FM-generated sub-task decompositions achieve high temporal, $\tau_{k}$, and semantic $\tau_{\zeta}$, similarity scores when guided with an in-context example, $\textbf{C}^{1S}$. This is because $\textbf{C}^{1S}$ directly steers the FM to attend to important transitional steps and to synthesize descriptors using terminology similar to the example.
For example, GPT-4V with textual input, $\mathcal{D}^{kxu}$, significantly outperforms both its zero-shot counterpart $\textbf{C}^{0S}$ and human annotations. The ablation study further reveals that including an in-context example not only improves alignment metrics but also curbs overly fine-grained predictions, while using visual data alone $\mathcal{D}^{\mu}$ tends to reduce semantic fidelity. The very low scores are consistent for $\hat{\mathcal{S}}$ that contain $N_{\hat{\mathcal{S}}}\gg N_{\mathcal{S}}$, indicating too fine of a granularity. For example, Video-LLaVa produced $N_{\hat{\mathcal{S}}}\approx K_{\mathcal{S}}$, implying a sub-task at every $k$, which is not realistic.
In practice, our FM-based annotation framework enables dataset curators to validate only a small set of FM-generated annotations against hand-annotated trajectories, dramatically reducing annotation time and cost. This approach facilitates the scalable production of high-quality, language-supervised datasets.
LLM (Context, Data) Door Lift PickPlace Stack
GPT-4V $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ 50 50 50 49
GPT-4V $(\textbf{C}^{0S},\mathcal{D}^{kxu})$ 50 50 50 50
GPT-4V $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ 50 49 50 50
GPT-4V $(\textbf{C}^{0S},\mathcal{D}^{\mu})$ 50 50 50 50
GPT-4V $(\textbf{C}^{1S},\mathcal{D})$ 50 50 50 49
GPT-4V $(\textbf{C}^{0S},\mathcal{D})$ 49 50 50 50
Gemini Pro $(\textbf{C}^{1S},\mathcal{D}^{kxu})$ 42 48 50 50
Gemini Pro $(\textbf{C}^{0S},\mathcal{D}^{kxu})$ 47 43 47 49
Gemini Pro V $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ 23 50 29 19
Gemini Pro V $(\textbf{C}^{0S},\mathcal{D}^{\mu})$ 47 34 44 38
Video-LLaVA $(\textbf{C}^{1S},\mathcal{D}^{\mu})$ 50 50 50 12
Video-LLaVA $(\textbf{C}^{0S},\mathcal{D}^{\mu})$ 50 50 50 50
TABLE III: The number of correctly formatted $\hat{\mathcal{S}}$, according to Definition 3, for each 50 trajectory batch. The number in each cell is the corresponding sample size for mean and standard deviation statistics of each $\tau_{k}$, $\tau_{\zeta}$ pair in Table II.
@@ VI Conclusion
We present a method to send multi-modal input data including step-by-step state, action and corresponding images to FMs including LLMs and VLMs. We steer an FM for use as an analytics engine to digest raw trajectory data in the prompt and query the FM to decompose the data into distinct sub-tasks. As of this writing, the cost of offloading and requesting data is \$0.03 to \$0.15 depending on the trajectory length. For a human annotator to be more cost-effective, they would need to annotate a trajectory in about 15-30 seconds, which is not consistent with our collected human annotation data of 2-3 minutes per trajectory. We see this analysis and capability of using FMs for robotic’s data labeling as cost-effective and data efficient for the broader research problem of robot learning at scale.
The results are shown with a ground-truth $\mathcal{S}$ crafted by a customized FSM. There may be other $\mathcal{S}$ with different granularity of sub-tasks, resulting in a different number of sub-tasks, $N_{\mathcal{S}}$, and alternative descriptors. Our results are valid because the proposed annotation pipeline and evaluation method will be set up by developers who want FM predictions, $\hat{\mathcal{S}}$, and similar to $\mathcal{S}$ of their own.
The IOU calculation is a strict temporal calculation, resulting in strict weightings. If this causes challenges for future developers, an alternative “soft” IOU may be utilized where an intersection may be greater than zero for temporal ranges that are near each other, but do not necessarily intersect. In our algorithm, shorter sub-tasks are assigned smaller weights towards $\tau_{k}$ and $\tau_{\zeta}$. In practice, a short sub-task may be more important to a developer than longer sub-tasks. Users can manually increase the relative weighting of important sub-tasks to give more contribution towards the final scores. While we use the USE encoder to calculate $\tau_{\zeta}$, our framework is designed to be modular, allowing future use of any pre-trained encoder.
The presented robotics manipulator environments have structured and repeatable $\mathcal{S}$, which allow us to present quantitative results. The presented simulated environments offer perfect robot and non-robot pose information, which may not be assumed in the real-world. We believe our experiments using only $\mathcal{D}^{\mu}$ input, which does not contain pose information, closely represents real-world conditions.
Experimentation can be extended in several manners: additional visual data inputs per trajectory through multiple camera perspectives, physical robots with noisy positional data, non-manipulator environments, etc. However, we believe the presented experiments and quantitative results are sufficient to showcase our proposed method.